# Transforming H&E images into IHC: A Variance-Penalized GAN for Precision Oncology

### v1

@@ Abstract
The overexpression of the human epidermal growth factor receptor 2 (HER2) in breast cells is a key driver of HER2-positive breast cancer, a highly aggressive subtype requiring precise diagnosis and targeted therapy. Immunohistochemistry (IHC) is the standard technique for HER2 assessment but is costly, labor-intensive, and highly dependent on antibody selection. In contrast, hematoxylin and eosin (H&E) staining, a routine histopathological procedure, offers broader accessibility but lacks HER2 specificity. This study proposes an advanced deep learning-based image translation framework to generate high-fidelity IHC images from H&E-stained tissue samples, enabling cost-effective and scalable HER2 assessment. By modifying the loss function of pyramid pix2pix, we mitigate mode collapse, a fundamental limitation in generative adversarial networks (GANs), and introduce a novel variance-based penalty that enforces structural diversity in generated images. Our model particularly excels in translating HER2-positive (IHC 3+) images, which have remained challenging for existing methods due to their complex morphological variations. Extensive evaluations on the BCI histopathological dataset demonstrate that our model surpasses state-of-the-art methods in terms of peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and Fréchet Inception Distance (FID), particularly in accurately translating HER2-positive (IHC 3+) images. Beyond medical imaging, our model exhibits superior performance in general image-to-image translation tasks, showcasing its potential across multiple domains. This work marks a significant step toward AI-driven precision oncology, offering a reliable and efficient alternative to traditional HER2 diagnostics.
Index Terms:
@@
I Introduction
Breast cancer remains the most prevalent form of cancer worldwide [1]. Among its various subtypes, approximately 20% of the cases are classified as HER2-positive breast cancer, characterized by overexpression of the human epidermal growth factor receptor 2 (HER2) [2]. This subtype is particularly aggressive and requires targeted therapies, such as Herceptin, to inhibit HER2 signaling. However, accurate HER2 classification is critical, as misdiagnosis can lead to inappropriate treatment strategies and higher recurrence rates [3].
The gold standard for HER2 detection is fluorescence in situ hybridization (FISH), a molecular technique that assesses gene amplification [4]. However, FISH is costly, time consuming and not widely available in all clinical laboratories [5]. Immunohistochemistry (IHC) provides a more accessible alternative, relying on antibody-antigen interactions to detect HER2 expression [6, 7]. IHC results are categorized into four levels: 0, 1+, 2+, and 3+, as illustrated in Fig. 1. Although IHC 0 and 1+ are HER2-negative, IHC 3+ is strongly HER2-positive and indicates eligibility for HER2-targeted therapies. However, IHC 2+ remains ambiguous, necessitating additional FISH testing [2]. Despite its utility, IHC still requires specialized reagents, expensive equipment, and experienced pathologists for interpretation, introducing subjectivity in diagnosis [8, 9].
@@ I Introduction
Figure 1: Different types of IHC images based on levels of HER2 expression in the BCI dataset [10]. The higher the HER2 expression, the greater the corresponding label.
Hematoxylin and eosin (H&E) staining is the standard histopathological technique used for initial cancer screening. Unlike IHC, H&E is cost-effective, rapid, and independent of antibody selection, making it widely used for cancer diagnosis. H&E stains nuclei blue to dark purple and cytoplasm in various shades of pink and red, providing high-contrast images for structural analysis. Fig. 2 illustrates the three primary techniques used in HER2 detection: H&E, IHC, and FISH.
@@
Figure 2: Different techniques for HER2 detection. These patches are from whole-slide images of biopsied tissue samples treated using H&E (a), IHC (b), and FISH (c).
Recent advances in deep learning have demonstrated that H&E images alone can predict HER2 status without IHC [11]. Studies [12, 13, 14, 15, 16] have used convolutional neural networks (CNNs) and transformers to classify HER2 expression directly from H&E images. However, a direct H&E-based classification lacks interpretability, which limits its adoption in clinical practice. Instead, a more promising approach is to translate H&E images into IHC-like images, allowing human pathologist validation and dataset augmentation.
Liu et al. [10] pioneered this approach using a GAN-based pyramid pix2pix model, training it on their BCI dataset. While pyramid pix2pix improved image translation, it struggled to generate realistic IHC 3+ images, suffering from mode collapse, where the model fails to capture diverse HER2-positive morphologies. This limitation undermines the clinical utility of the generated IHC images, particularly for HER2-positive cases where a precise classification is required.
@@
To achieve these objectives, we introduce a variance penalty term into the loss function of the pyramid pix2pix model, enforcing structural diversity and fidelity in IHC image generation. Our model is rigorously evaluated using the BCI dataset for HER2 classification and LLVIP dataset for general image translation, demonstrating superior performance over existing approaches [17].
@@
The remainder of this paper is structured as follows: Section II reviews existing research on breast cancer detection and image translation techniques. Section III outlines our proposed methodology. Section IV describes the experimental setup including the analysis of the dataset and the discussion of evaluation metrics we have used. Section V presents the comparative results, and Section VI concludes the study with insights on future research directions.
II Literature Review
@@
Breast cancer is among the most prevalent malignancies worldwide, requiring advanced techniques for early detection and classification. Significant research efforts have been devoted to automating breast cancer diagnosis using a combination of image processing, machine learning, and deep learning techniques. Traditional image processing methods have focused on nuclei segmentation and feature extraction to differentiate between malignant and benign tumors [18]. With the advent of machine learning, algorithms such as support vector machines (SVMs) and random forests have demonstrated high efficacy in breast cancer classification [19, 20, 21]. More recently, deep learning architectures, particularly convolutional neural networks (CNNs), have significantly enhanced histopathological image classification by automatically learning hierarchical representations [22, 23, 24, 25].
@@
The recent emergence of Vision Transformers (ViTs) [26] and their variants [27, 28] has further revolutionized the field. Unlike CNNs, which rely on localized receptive fields, ViTs utilize self-attention mechanisms to model long-range dependencies in images, leading to improved feature extraction and classification. Several studies have successfully deployed ViTs for breast cancer detection [29, 30, 31, 32, 33], demonstrating superior performance over traditional CNNs.
@@
An important aspect of automated breast cancer diagnosis is the availability of high-quality datasets. The BreakHis dataset [34] is widely used in breast cancer classification, consisting of 9,109 microscopic images of benign and malignant tumors obtained at different magnifications. Additionally, the BreCaHAD (Breast Cancer Histopathological Annotation and Diagnosis) dataset [35] and the BACH (Breast Cancer Histology Challenge) dataset [36] provide rich sources of labeled histopathological images for benchmarking various computational models.
@@
Beyond morphological classification, biomarker analysis plays a pivotal role in determining treatment strategies for breast cancer patients. Immunohistochemistry (IHC) is a widely used method to assess biomarkers such as estrogen receptors (ER), progesterone receptors (PR), and human epidermal growth factor receptor 2 (HER2), which guide targeted therapies. Recent studies [11, 13, 12] have demonstrated that machine learning algorithms can predict biomarker expression directly from hematoxylin and eosin (H&E)-stained images, highlighting a strong correlation between morphological structures in H&E images and biochemical properties observed in IHC images. This has motivated researchers to explore the translation of H&E images into IHC images as a means to facilitate biomarker assessment while reducing costs and dependency on specialized laboratory techniques.
@@
Image-to-image translation is a critical area of research in computational pathology, particularly for the transformation of H&E-stained images into IHC images. Liu et al. [10] pioneered this effort by introducing a supervised deep learning framework for H&E to IHC image translation. Broadly, image translation methods can be categorized into unsupervised and supervised approaches. Unsupervised translation models, such as CycleGAN [37], DualGAN [38], and UNIT (Unsupervised Image-to-Image Translation Network) [39], leverage cycle consistency losses to map images between domains without paired training data. However, these approaches often struggle with preserving fine-grained histopathological details. In contrast, supervised models utilize paired datasets to learn a direct mapping between input and target images, resulting in more precise translations.
Among the supervised image translation models, the pix2pix framework [40] introduced conditional generative adversarial networks (cGANs) for image-to-image translation. pix2pixHD [41] extended this approach to generate high-resolution images using coarse-to-fine generators and multi-scale discriminators. Building on these advancements, Liu et al. proposed pyramid pix2pix, an enhanced version of pix2pix tailored for translating H&E images into IHC images [10]. Their key contribution was the creation of the BCI dataset, a large-scale paired dataset of H&E and IHC images. While pyramid pix2pix improved translation quality for low HER2 expression levels (IHC 0/1+/2+), it failed to generate realistic IHC 3+ images, suffering from mode collapse, a common limitation in generative adversarial networks (GANs).
@@
Addressing mode collapse remains a critical challenge in GAN-based image translation. Several techniques have been proposed to mitigate this issue, including Wasserstein GANs (WGANs) [42], spectral normalization [43], mode regularization [44], and self-attention mechanisms [45]. However, none of these strategies have been explicitly applied to the translation of H&E images into IHC images. This gap underscores the need for novel loss functions that penalize mode collapse while maintaining fidelity and structural diversity in generated IHC images.
@@
Given these limitations, our study introduces an enhanced loss function that explicitly penalizes low-diversity image generation by incorporating a variance-based regularization term. By addressing mode collapse, our approach improves the quality and diversity of IHC 3+ image generation, thereby advancing computer-aided HER2 breast cancer diagnostics. The proposed methodology is not only specific to histopathological image translation but also demonstrates promising results for generic image-to-image translation tasks, extending its applicability beyond biomedical imaging.
III Proposed Methodology
@@
The transformation of hematoxylin and eosin (H&E) stained images into Immunohistochemistry (IHC) stained images is a fundamental challenge in digital pathology. Accurate translation between these imaging modalities is essential for both diagnostic and prognostic applications. However, existing generative models, such as pyramid pix2pix [10], often suffer from mode collapse, leading to reduced variability in generated images and suboptimal generalization across various histopathological samples, especially in translating HER2-positive (IHC 3+) images.
@@
To overcome these challenges, we propose an enhancement to the loss function of pyramid pix2pix framework by introducing a variance difference regularization term. The general framework of our approach is illustrated in Fig. 3. The proposed modification encourages diversity in generated images, while maintaining fidelity to the original IHC image distributions. By penalizing lower variance in generated images relative to the input images, the model’s learning is driven to generate more diverse and realistic representations, thereby reducing mode collapse and improving translation accuracy.
@@
Figure 3: The framework of the proposed architecture for generating high-fidelty IHC images from H&E images. The relative variance between input images and generated images in a batch is incorporated into the overall loss of pyramid pix2pix [10].
III-A Pyramid pix2pix Framework
@@
$f$: Real Images Tensor $n\times C\times H\times W$ $g$: Generated Images Tensor $n\times C\times H\times W$ $\sigma^{2}(f)$: Variance for each pixel (Real Images) $\sigma^{2}(g)$: Variance for each pixel (Generated Images) Average Variance for Real Images $\bar{\sigma^{2}(f)}$ Average Variance for Generated Images $\bar{\sigma^{2}(g)}$ $g=|\bar{\sigma^{2}(f)}-\bar{\sigma^{2}(g)}|$: Absolute Difference of Averages Real and Generated Image TensorsVariance CalculationsFinal Output: Absolute Difference
Figure 4: Tensor operations and variance calculations for real and generated images.
III-B Variance-Based Loss for Mode Collapse Mitigation
@@
First, the mean pixel value at each channel and spatial location is calculated as:
$\mu(c,i,j)=\frac{1}{n}\sum_{k=1}^{n}{f}_{k,c,i,j}$ (6)
The variance at each position across the batch is then computed as:
$\sigma^{2}(c,i,j)=\frac{1}{n}\sum_{k=1}^{n}\left(f_{k,c,i,j}-\mu(c,i,j)\right)^{2}$ (7)
Averaging over all pixels and channels, we obtain the overall variance:
$\overline{\sigma^{2}}=\frac{1}{C\cdot H\cdot W}\sum_{c=1}^{C}\sum_{i=1}^{H}\sum_{j=1}^{W}\sigma^{2}(c,i,j)$ (8)
Let $\overline{\sigma^{2}(f)}$ and $\overline{\sigma^{2}(G(f))}$ denote the variance of real and generated images, respectively. The variance difference is then defined as:
$\Delta\sigma^{2}=\left|\overline{\sigma^{2}(f)}-\overline{\sigma^{2}(G(f))}\right|$ (9)
This variance regularization term ensures that the generator maintains diversity in generated outputs. Fig. 4 illustrates the tensor operations involved in computing variance.
III-C Final Objective Function

### v2

# Transforming Hematoxylin-Eosin Images into Immunohistochemistry Images: A Variance-Penalized GAN for Precision Oncology

@@ Abstract
The overexpression of the human epidermal growth factor receptor 2 (HER2) in breast cells is a key driver of HER2-positive breast cancer, a highly aggressive subtype requiring precise diagnosis and targeted therapy. Immunohistochemistry (IHC) is the standard technique for HER2 assessment but is costly, labor-intensive, and highly dependent on antibody selection. In contrast, hematoxylin and eosin (H&E) staining, a routine histopathological procedure, offers broader accessibility but lacks HER2 specificity. This study proposes an advanced deep learning-based image translation framework to generate high-fidelity IHC images from H&E-stained tissue samples, enabling cost-effective and scalable HER2 assessment. By modifying the loss function of pyramid pix2pix, we mitigate mode collapse, a fundamental limitation in generative adversarial networks (GANs), and introduce a novel variance-based penalty that enforces structural diversity in generated images. Our model particularly excels in translating HER2-positive (IHC 3+) images, which have remained challenging for existing methods. Quantitative evaluations on the overall BCI dataset reveal that our approach outperforms baseline models, achieving a peak signal-to-noise ratio (PSNR) of 22.16, a structural similarity index (SSIM) of 0.47, and a Fréchet Inception Distance (FID) of 346.37. In comparison, the pyramid pix2pix baseline attained PSNR 21.15, SSIM 0.43, and FID 516.75, while the standard pix2pix model yielded PSNR 20.74, SSIM 0.44, and FID 472.6. These results affirm the superior fidelity and realism of our generated IHC images. Beyond medical imaging, our model exhibits superior performance in general image-to-image translation tasks, showcasing its potential across multiple domains. This work marks a significant step toward AI-driven precision oncology, offering a reliable and efficient alternative to traditional HER2 diagnostics.
[…]
1 Keywords:
@@
2 Introduction
Breast cancer remains the most prevalent form of cancer worldwide (lei2021global). Among its various subtypes, approximately 20% of the cases are classified as HER2-positive breast cancer, characterized by overexpression of the human epidermal growth factor receptor 2 (HER2) (Siegel2022). This subtype is particularly aggressive and requires targeted therapies, such as Herceptin, to inhibit HER2 signaling. However, accurate HER2 classification is critical, as misdiagnosis can lead to inappropriate treatment strategies and higher recurrence rates (slamon1987human).
The gold standard for HER2 detection is fluorescence in situ hybridization (FISH), a molecular technique that assesses gene amplification (prati2005histopathologic). However, FISH is costly, time consuming and not widely available in all clinical laboratories (ross2004targeted). Immunohistochemistry (IHC) provides a more accessible alternative, relying on antibody-antigen interactions to detect HER2 expression (thomson2001her, van2001assessment). IHC results are categorized into four levels: 0, 1+, 2+, and 3+, as illustrated in Figure 1. Although IHC 0 and 1+ are HER2-negative, IHC 3+ is strongly HER2-positive and indicates eligibility for HER2-targeted therapies. However, IHC 2+ remains ambiguous, necessitating additional FISH testing (Siegel2022). Despite its utility, IHC still requires specialized reagents, expensive equipment, and experienced pathologists for interpretation, introducing subjectivity in diagnosis (robb2015call, lin2014standardization).
@@ 2 Introduction
Figure 1: Different types of IHC images based on levels of HER2 expression in the BCI dataset (liu2022bci). The higher the HER2 expression, the greater the corresponding label.
Hematoxylin and eosin (H&E) staining is the standard histopathological technique used for initial cancer screening. Unlike IHC, H&E is cost-effective, rapid, and independent of antibody selection, making it widely used for cancer diagnosis. H&E stains nuclei blue to dark purple and cytoplasm in various shades of pink and red, providing high-contrast images for structural analysis. Figure 2 illustrates the three primary techniques used in HER2 detection: H&E, IHC, and FISH.
@@
Figure 2: Different techniques for HER2 detection.
Recent advances in deep learning have demonstrated that H&E images alone can predict HER2 status without IHC (shamai2019artificial). Studies (gamble2021determining, naik2020deep, bychkov2021deep, rawat2020deep, levy2020preliminary) have used convolutional neural networks (CNNs) and transformers to classify HER2 expression directly from H&E images. However, a direct H&E-based classification lacks interpretability, which limits its adoption in clinical practice. Instead, a more promising approach is to translate H&E images into IHC-like images, allowing human pathologist validation and dataset augmentation.
Liu et al. (liu2022bci) pioneered this approach using a GAN-based pyramid pix2pix model, training it on their BCI dataset. While pyramid pix2pix improved image translation, it struggled to generate realistic IHC 3+ images, suffering from mode collapse, where the model fails to capture diverse HER2-positive morphologies. This limitation undermines the clinical utility of the generated IHC images, particularly for HER2-positive cases where a precise classification is required.
@@
To achieve these objectives, we introduce a variance penalty term into the loss function of the pyramid pix2pix model, enforcing structural diversity and fidelity in IHC image generation. Our model is rigorously evaluated using the BCI dataset for HER2 classification and the facades dataset (isola2017image) for general image translation, demonstrating superior performance over existing approaches.
@@
The remainder of this paper is organized as follows. Section 2 reviews the related literature on breast cancer detection and image-to-image translation techniques. Section 3 presents the proposed methodology. Section 4 describes the experimental setup, including dataset analysis and the evaluation metrics employed. Sections 5 and 6 report and discuss the comparative results. Finally, Section 7 concludes the study and outlines potential directions for future research.
Source Code: The algorithmic implementation, developed in Python, is openly available for replication and further development at our GitHub repository: https://github.com/SaraRehmat/HE2IHC.
3 Literature Review
@@
Breast cancer is among the most prevalent malignancies worldwide, requiring advanced techniques for early detection and classification. Significant research efforts have been devoted to automating breast cancer diagnosis using a combination of image processing, machine learning, and deep learning techniques. Traditional image processing methods have focused on nuclei segmentation and feature extraction to differentiate between malignant and benign tumors (xu2014efficient). With the advent of machine learning, algorithms such as support vector machines (SVMs) and random forests have demonstrated high efficacy in breast cancer classification (wang2018support, kumar2020prediction, lahoura2021cloud). More recently, deep learning architectures, particularly convolutional neural networks (CNNs), have significantly enhanced histopathological image classification by automatically learning hierarchical representations (hirra2021breast, jiang2019breast, khan2019novel, xie2019deep).
@@
The recent emergence of Vision Transformers (ViTs) (dosovitskiy2020image) and their variants (touvron2021training, liu2021swin) has further revolutionized the field. Unlike CNNs, which rely on localized receptive fields, ViTs utilize self-attention mechanisms to model long-range dependencies in images, leading to improved feature extraction and classification. Several studies have successfully deployed ViTs for breast cancer detection (alotaibi2023vit, tummala2022breast, shao2021transmil, gul2022histopathological, wang2022semi), demonstrating superior performance over traditional CNNs.
@@
An important aspect of automated breast cancer diagnosis is the availability of high-quality datasets. The BreakHis dataset (spanhol2015dataset) is widely used in breast cancer classification, consisting of 9,109 microscopic images of benign and malignant tumors obtained at different magnifications. Additionally, the BreCaHAD (Breast Cancer Histopathological Annotation and Diagnosis) dataset (aksac2019brecahad) and the BACH (Breast Cancer Histology Challenge) dataset (aresta2019bach) provide rich sources of labeled histopathological images for benchmarking various computational models.
@@
Beyond morphological classification, biomarker analysis plays a pivotal role in determining treatment strategies for breast cancer patients. Immunohistochemistry (IHC) is a widely used method to assess biomarkers such as estrogen receptors (ER), progesterone receptors (PR), and human epidermal growth factor receptor 2 (HER2), which guide targeted therapies. Recent studies (shamai2019artificial, naik2020deep, gamble2021determining) have demonstrated that machine learning algorithms can predict biomarker expression directly from hematoxylin and eosin (H&E)-stained images, highlighting a strong correlation between morphological structures in H&E images and biochemical properties observed in IHC images. This has motivated researchers to explore the translation of H&E images into IHC images as a means to facilitate biomarker assessment while reducing costs and dependency on specialized laboratory techniques.
@@
Image-to-image translation is a critical area of research in computational pathology, particularly for the transformation of H&E-stained images into IHC images. Liu et al. (liu2022bci) pioneered this effort by introducing a supervised deep learning framework for H&E to IHC image translation. Broadly, image translation methods can be categorized into unsupervised and supervised approaches. Unsupervised translation models, such as CycleGAN (zhu2017unpaired), DualGAN (yi2017dualgan), and UNIT (Unsupervised Image-to-Image Translation Network) (liu2017unsupervised), leverage cycle consistency losses to map images between domains without paired training data. TC-CycleGAN (huang2023tc) proposed an improved CycleGAN-based unpaired translation framework to translate H&E images into IHC images. The method introduced self-attention modules in both the generator and discriminator to better capture long-range dependencies, and a Gabor-based texture loss to preserve fine texture details.However, this and other unsupervised translation methods often struggle with preserving fine-grained histopathological details. In contrast, supervised models utilize paired datasets to learn a direct mapping between input and target images, resulting in more precise translations.
Among the supervised image translation models, the pix2pix framework (isola2017image) introduced conditional generative adversarial networks (cGANs) for image-to-image translation. Pix2pixHD (wang2018high) extended this approach to generate high-resolution images using coarse-to-fine generators and multi-scale discriminators. Building on these advancements, Liu et al. proposed pyramid pix2pix, an enhanced version of pix2pix tailored for translating H&E images into IHC images (liu2022bci). Their key contribution was the creation of the BCI dataset, a large-scale paired dataset of H&E and IHC images. While pyramid pix2pix improved translation quality for low HER2 expression levels (IHC 0/1+/2+), it failed to generate realistic IHC 3+ images, suffering from mode collapse, a common limitation in generative adversarial networks (GANs). (Liu2023MGGANAM) suggested using multi-generator in their GAN-based architecture. The multi-generator has two U-shaped networks at different scales designed to capture low-frequency components like structure and global layout and high frequency components like texture and staining details. They also modified the loss function of pix2pix by adding a cross-entropy regularization term. But their method also struggled to accurately translate IHC 3+ samples and identify dark regions with high HER2 expression. IHC-GAN (saad2025automatic) used a similar dual-scale generator. The framework first employed a MobileNetV3 classifier to predict the HER2 expression level (0, 1+, 2+, or 3+) from the input H&E image and extract discriminative feature embeddings. The features were then incorporated into a dual-scale generator through Adaptive Instance Normalization (AdaIN), enabling the generator to modulate staining style according to the predicted HER2 stage. The generator consisted of two coupled encoder–decoder branches with inverted residual blocks to reduce parameters. The model used a composite loss that combined adversarial loss, L1 reconstruction loss, perceptual (VGG-based) loss, and feature-matching loss to balance visual fidelity and structural consistency. The pipeline combined a classifier, dual-scale generator, AdaIN conditioning, and multi-scale discriminator, resulting in a structurally complex system that could lead to overfitting in limited medical datasets.
@@
Addressing mode collapse remains a critical challenge in GAN-based image translation. Several techniques have been proposed to mitigate this issue, including Wasserstein GANs (WGANs) (adler2018banach), spectral normalization (miyato2018spectral), mode regularization (che2016mode), and self-attention mechanisms (zhang2019self). However, none of these strategies have been explicitly applied to the translation of H&E images into IHC images. This gap underscores the need for novel loss functions that penalize mode collapse while maintaining fidelity and structural diversity in generated IHC images.
@@
Given these limitations, our study introduces an enhanced loss function that explicitly penalizes low-diversity image generation by incorporating a variance-based regularization term. By addressing mode collapse, our approach improves the quality and diversity of IHC image generation, thereby advancing computer-aided HER2 breast cancer diagnostics. The proposed methodology is not only specific to histopathological image translation but also demonstrates promising results for generic image-to-image translation tasks, extending its applicability beyond biomedical imaging.
4 Proposed Methodology
@@
The transformation of hematoxylin and eosin (H&E) stained images into Immunohistochemistry (IHC) stained images is a fundamental challenge in digital pathology. Accurate translation between these imaging modalities is essential for both diagnostic and prognostic applications. However, existing generative models, such as pyramid pix2pix (liu2022bci), often suffer from mode collapse, leading to reduced variability in generated images and suboptimal generalization across various histopathological samples, especially in translating HER2-positive (IHC 3+) images.
@@
To overcome these challenges, we propose an enhancement to the loss function of pyramid pix2pix framework by introducing a variance difference regularization term. The general framework of our approach is illustrated in Figure 3. The proposed modification encourages diversity in generated images, while maintaining fidelity to the original IHC image distributions. By penalizing lower variance in generated images relative to the input images, the model’s learning is driven to generate more diverse and realistic representations, thereby reducing mode collapse and improving translation accuracy.
@@
Figure 3: The framework of the proposed architecture for generating high-fidelty IHC images from H&E images. The relative variance between input images and generated images in a batch is incorporated into the overall loss of pyramid pix2pix (liu2022bci).
4.1 Pyramid pix2pix Framework
@@
4.2 Variance-Based Loss for Mode Collapse Mitigation
@@
First, the mean pixel value at each channel and spatial location is computed across the batch as:
$\mu_{f}(c,i,j)=\frac{1}{n}\sum_{k=1}^{n}f_{k,c,i,j},\qquad\mu_{g}(c,i,j)=\frac{1}{n}\sum_{k=1}^{n}G(f)_{k,c,i,j}$ (6)
The batch-wise variance at each channel and spatial position is then:
$\sigma_{f}^{2}(c,i,j)=\frac{1}{n}\sum_{k=1}^{n}\left(f_{k,c,i,j}-\mu_{f}(c,i,j)\right)^{2}$ (7)
$\sigma_{g}^{2}(c,i,j)\frac{1}{n}\sum_{k=1}^{n}\left(G(f)_{k,c,i,j}-\mu_{g}(c,i,j)\right)^{2}$ (8)
Next, the absolute variance difference is computed at each pixel location:
$\Delta\sigma^{2}(c,i,j)=\left|\sigma_{f}^{2}(c,i,j)-\sigma_{g}^{2}(c,i,j)\right|$ (9)
Finally, the proposed variance regularization loss is obtained by averaging these absolute differences over all channels and spatial locations:
$\mathcal{L}_{var}=\frac{1}{C\cdot H\cdot W}\sum_{c=1}^{C}\sum_{i=1}^{H}\sum_{j=1}^{W}\Delta\sigma^{2}(c,i,j)$ (10)
This variance regularization term ensures that the generator maintains diversity in generated outputs. Figure 4 illustrates the tensor operations involved in computing variance.
4.3 Final Objective Function